# Modulo Sampling: Performance Guarantees in The Presence of Quantization

### v1

@@ III-B MSE Guarantee for Modulo ADC Without 1-Bit Folding Information
(from [37, Theorem 2]) Consider the observation signal $\mathbf{y}=\mathbf{A}\mathbf{x}+\mathbf{e}$. Suppose that $\|\mathbf{A}^{\mathrm{H}}\mathbf{e}\|\leq\epsilon_{0}$ and matrix $\mathbf{A}$ satisfies
@@
where $\delta_{L}$ is the restricted isometry property (RIP) constant order $L$ of matrix $\mathbf{A}$. Then, the OMP with stopping rule $\|\mathbf{A}^{\mathrm{H}}\mathbf{e}\|\leq\epsilon_{0}$ will exactly recover the support $\Omega$ of an $L$-sparse signal $\mathbf{x}$ from the observation signal $\mathbf{y}$ if the minimum magnitude of nonzero elements of $\mathbf{x}$ satisfies
$\min_{i\in\Omega}|\mathbf{x}_{i}|>\frac{\left(\sqrt{1+\delta_{L+1}}+1\right)\sqrt{L}\epsilon_{0}}{1-(\sqrt{L}+1)\delta_{L+1}}.$ (9)
Using Theorem 2, we identify a sufficient condition on the quantizer resolution $b$ to ensure that OMP will produce $\hat{c}[n]=c[n]$. With pefect $\hat{c}[n]$, we can apply Theorem 1 to obtain recovery guarantees for the proposed algorithm in Section II-C.
@@
Set $\lambda^{\prime}=\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}$ and let
@@
Suppose matrix $\mathbf{V}$ satisfies the RIP of order $L_{0}$ with an RIP constant denoted as $\delta_{L}$ and that $\delta_{L_{0}}<\frac{1}{\sqrt{L_{0}}+1}$. Let $\eta=\|\mathbf{V}^{\mathrm{H}}\mathbf{V}\|_{\infty}$. If $\mathrm{OF}>3$ and
@@
then $c[n]$ can be perfectly recovered using OMP with stopping rule $\|\mathbf{V}^{\mathrm{H}}\mathbf{r}_{t}\|_{\infty}\leq\eta\frac{6\lambda^{\prime}}{2^{b}-2}$. The MSE of the OMP-based reconstruction procedure for the modulo ADC system without the 1-bit folding information signal is
$\mathrm{MSE}=$ $\frac{1}{\mathrm{OF}(2^{b}-2)^{2}}\cdot\left(\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}\right)^{2}.$ (12)
@@
The main difference between Theorem 3 and Theorem 1 is the quantization bit requirement needed to obtain the performance guarantees. The argument inside the logarithm in (11) can be shown to be greater than 1. Thus, the sufficient condition on the quantization bits becomes more stringent in the absence of 1-bit folding information $c[n]$. One major limitation of the bound is that it depends on the RIP constant $\delta_{L_{0}+1}$. Testing whether a sensing matrix satisfies RIP is an NP-hard problem [38]. We also note that a different sparse recovery algorithm other than OMP can be used to recover $c[n]$. For instance, the LASSO-B2R2 [15] uses the iterative soft-thresholding algorithm (ISTA) for sparse recovery.
@@ IV Case Study: Simultaneous Acquisition of Weak and Strong Signals
We now investigate the case of simultaneous acquisition of weak and strong signals that occupy different frequency bands, as shown in Figure 3. The modulo ADC block is based on Fig. 2a and the digital channelizer implements a digital bandpass filter that can be tuned to the desired channel. This case is particularly interesting due to the conflicting requirements to capture the two signals. Setting the dynamic range too high to capture the strong signal can bury the weak signal in quantization noise. Meanwhile, setting the dynamic range too low could distort the strong signal. A recent work on USF [28] considers the issue of weak and strong components. However, their mathematical formulation of the problem only assumes a two-tone input signal with different amplitudes. Moreover, performance improvements of modulo ADC over conventional ADCs are only presented via numerical simulations. To the best of our knowledge, a theoretical analysis of the simultaneous acquisition of weak and strong signals using modulo ADC has not been conducted. Our goal in this section is to provide a more general mathematical formulation of the problem and a rigorous analysis to justify the advantage of modulo ADC over conventional ADCs in this case.
@@
where $f_{1}(t)$ and $f_{2}(t)$ are bandpass signals. Denote $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ the frequency supports of $f_{1}(t)$ and $f_{2}(t)$, respectively. We impose $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\varnothing$ and $\mathcal{F}_{1}\cup\mathcal{F}_{2}\in[-\omega_{m}/2,+\omega_{m}/2]$, i.e., $f_{1}(t)$ and $f_{2}(t)$ have different frequency supports and they both reside inside $[-\omega_{m}/2,+\omega_{m}/2]$. The bandwidths of $f_{1}(t)$ and $f_{2}(t)$ are $\omega_{1}$ and $\omega_{2}$, respectively. Without loss of generality, we set $\|f_{1}(t)\|_{2}^{2}=1$ and $\|f_{2}(t)\|_{2}^{2}=1$. The simultaneous acquisition of strong and weak signals is modeled by setting $\alpha_{1}\gg\alpha_{2}$. Since $f(t)$ is a bandlimited signal with energy $\alpha_{1}^{2}+\alpha_{2}^{2}$, we have $\|f(t)\|_{\infty}\leq\sqrt{\frac{\omega_{m}(\alpha_{1}^{2}+\alpha_{2}^{2})}{\pi}}$ by [39].
@@
At this point, one might consider separating the two signal components in the analog domain using a tunable bandpass filter. This approach suppresses the strong signal component before it reaches the ADC. However, such an approach is costly to implement in analog. Moreover, analog filters exhibit severe group delay near the band edges, causing signal distortion. In contrast, a tunable digital filter can be implemented by varying the filter coefficients in the registers. Furthermore, achieving a linear phase response is straightforward in digital filters by imposing symmetry on the filter coefficients [40].
@@
Suppose $b>3$ and $\mathrm{OF}>3$. Then the NMSE incurred when recovering the $i$-th component under the modified reconstruction procedure is
@@ V Numerical Results
In this section, we perform the following simulations to substantiate the proposed theory. The aim is to show that the modulo ADC surpasses the conventional ADC lacking the modulo operator in terms of quantization noise reduction. This is accomplished by examining the variation of quantization noise power, represented by the MSE, as the OF value varies.
Simulation-1: In this simulation, we consider the modulo ADC system with side information as depicted in Fig. 2a. Specifically, we analyze the variation of MSE, as described in Theorem 1 and (7), as the OF value varies.
@@
where $\text{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ for any $x\neq 0$. This signal entails a maximal frequency component of $f_{m}=50$ Hz, and a temporal duration of $T=1$ second. Random coefficients $A_{i}$ are drawn uniformly from the range [1, 1]11[-1,\ 1][ - 1 , 1 ]. While time offsets $\tau_{i}$ are chosen randomly within the interval $\left(\frac{T}{4},\ \frac{3T}{4}\right)$, this is to ensure that the residual signal remains zero outside the interval $(0,\ T)$. Additionally, we normalize $f(t)$ such that $\|f(t)\|_{\infty}=1$. Consequently, the modulo threshold is $\lambda^{\prime}=\frac{1}{\mathrm{OF}-2}$. We consider the number of bits $b$ equal to $4$ and sweep the OF values from $3$ to $15$. The dither signal and uniform quantizer are configured according to these parameters. For each OF value, we compute the MSE between true samples, $f[n]$, and the estimated samples, $\hat{f}[n]$. The experiment is repeated for $20000$ i.i.d. noise realizations and the errors are averaged over all realizations.
@@
Fig. 4 depicts the results obtained with this setting. It is evident from the figure that the MSE of the modulo ADC decays much faster than that of the conventional ADC without the modulo operator. Therefore, from Fig. 4, we conclude that the quantization noise suppression of modulo ADC is better than that of conventional ADC. Moreover, it is worth noting that for $\text{OF}=3$, $\lambda^{\prime}=\|f(t)\|_{\infty}$, and the MSE values of both modulo ADC and conventional ADC are equal. Fig. 4 also depicts theoretical MSE values obtained using Theorem 1 and (7). Note that both theoretical predictions match the simulated results. This justifies the accuracy of our analytical results and further emphasizes the fact that the MSE decays with the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ and $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ for modulo ADC and conventional ADC, respectively. As a final remark, the use of the extra 1-bit side information in Fig. 2a only incurs one additional OR gate to the setup [16]. This penalty is fixed regardless of $b$. From a power efficiency viewpoint, this additional OR gate has minimal impact on the power consumption of modulo ADC.
Simulation-2: In this simulation, we examine the modulo ADC system without the additional 1-bit side information. Specifically, we analyze the MSE variation, as detailed in Theorem 3 and (7), across different OF values. The input signal settings align with those of Simulation-1, featuring $b=4$, OF ranging from $4$ to $15$, an OMP stopping criteria parameter of $\eta=0.8$, and $20000$ i.i.d noise realizations. Fig. 5 shows the variation of MSE, between true samples and estimated samples, for this setting as OF varies. The figure clearly demonstrates that modulo ADCs with and without extra-bit side information exhibit superior quantization noise suppression compared to the conventional ADC. Furthermore, our theoretical predictions align well with the simulated results; highlighting that the MSE diminishes at rates of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ and $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ for the modulo ADC and conventional ADC, respectively.
@@
Simulation-4:
@@
Figure 8: Numerical MSE vs. OF result for modulo ADC (with ) with and without extra 1-bit side information, and conventional ADC.
@@
Both previous simulations use the dither quantization framework. Here, we demonstrate that even without a dither signal (i.e., using the conventional quantization framework), the modulo ADC combined with the proposed recovery algorithms outperforms the classical ADC. For this analysis, we consider an input signal with the same settings as Simulation-1: $b=4$, an oversampling factor (OF) ranging from $4$ to $11$, an OMP stopping criterion parameter of $\eta=0.8$, a modulo threshold $\lambda=\frac{1}{\mathrm{OF}-2}$, and $5000$ i.i.d. noise realizations. Fig. 8 illustrates the variation in MSE across different OF values for both modulo and classical ADCs. The results show that the modulo ADC, with or without 1-bit information, consistently outperforms the classical ADC.
VI Summary

### v2

@@ III-B MSE Guarantee for Modulo ADC Without 1-Bit Folding Information
(from [49, Theorem 2]) Consider the observation signal $\mathbf{y}=\mathbf{A}\mathbf{x}+\mathbf{e}$. Suppose that $\|\mathbf{A}^{\mathrm{H}}\mathbf{e}\|_{\infty}\leq\epsilon_{0}$ and matrix $\mathbf{A}$ satisfies
@@
where $\delta_{L}$ is the restricted isometry property (RIP) constant order $L$ of matrix $\mathbf{A}$. Then, the OMP with stopping rule $\|\mathbf{A}^{\mathrm{H}}\mathbf{e}\|_{\infty}\leq\epsilon_{0}$ will exactly recover the support $\Omega$ of an $L$-sparse signal $\mathbf{x}$ from the observation signal $\mathbf{y}$ if the minimum magnitude of nonzero elements of $\mathbf{x}$ satisfies
$\min_{i\in\Omega}|\mathbf{x}_{i}|>\frac{\left(\sqrt{1+\delta_{L+1}}+1\right)\sqrt{L}\cdot\epsilon_{0}}{1-(\sqrt{L}+1)\delta_{L+1}}.$ (14)
Using Theorem 2, we identify a sufficient condition on the quantizer resolution $b$ to ensure that OMP will produce $\hat{c}[n]=c[n]$. With perfect $\hat{c}[n]$, we can apply Theorem 1 to obtain recovery guarantees for the proposed algorithm in Section II-C.
@@
Set $\lambda^{\prime}=\frac{c\cdot\|f(t)\|_{\infty}}{\mathrm{OF}-2}$ for some positive constant $c\in[1,\mathrm{OF}-2]$ and choose $n_{0}$, $n_{1}$ such that truncation error is negligible. Let
@@
Suppose matrix $\mathbf{V}$ satisfies the RIP of order $L_{0}$ with a RIP constant denoted $\delta_{L_{0}}$ and that $\delta_{L_{0}+1}<\frac{1}{\sqrt{L_{0}}+1}$. Let $\eta=\|\mathbf{V}^{\mathrm{H}}\mathbf{V}\|_{\infty}$. If $\mathrm{OF}>3$ and
@@
then $c[n]$ can be perfectly recovered using OMP with stopping rule $\|\mathbf{V}^{\mathrm{H}}\mathbf{r}_{t}\|_{\infty}\leq\eta\frac{6\lambda^{\prime}}{2^{b}-2}$. The $\mathrm{MSE}_{r}$ of the OMP-based reconstruction procedure for the modulo ADC system without the 1-bit folding information signal is
$\mathrm{MSE}_{r}=$ $\frac{c^{2}}{\mathrm{OF}(2^{b}-2)^{2}}\cdot\left(\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}\right)^{2}.$ (17)
@@
The intuition of the proof is to identify that the smallest non-zero magnitude of $\Delta z[n]$ is $\lambda^{\prime}$, and that $\epsilon_{0}=\eta\frac{6\lambda^{\prime}}{2^{b}-2}$ in (14). The resulting equation enabled us to derive a sufficient condition for the quantization bit.
@@
The main difference between Theorem 3 and Theorem 1 is the sufficient condition on the quantization bits to attain the MSE guarantees. The argument inside the logarithm in (16) can be shown to be greater than 1. Thus, the sufficient condition on the quantization bits becomes more stringent in the absence of 1-bit folding information $c[n]$. One major limitation of the bound is that it depends on the RIP constant $\delta_{L_{0}+1}$. Testing whether a sensing matrix satisfies RIP is an NP-hard problem [50]. Nevertheless, the sensing matrix $\mathbf{V}_{\mathcal{S}}$ used in our analysis closely resembles a partial DFT matrix. Partial DFT matrices are a class of structured random matrices that are widely studied in compressed sensing and are known to satisfy the RIP with high probability when the number of measurements is sufficiently large relative to the signal sparsity level [51]. This theoretical guarantee provides some intuitive justification of the RIP-based analysis in Theorem 3, even though the RIP constant cannot be evaluated explicitly in practice. Finally, we note that there are other sparse recovery algorithms that can be used to recover $c[n]$. For instance, the LASSO-B2R2 [20] uses the iterative soft-thresholding algorithm (ISTA) for sparse recovery. OMP is chosen in this work because of its established theoretical guarantees for perfect support recovery under the bounded noise setting, i.e., Theorem 2.
Remark: The above analysis provides a sufficient condition on OF and $b$ to ensure that the modulo ADC achieves lower MSE than a conventional ADC. That said, some recent works [47, 52] have explored the use of 1-bit quantizers in both conventional [52] and modulo ADCs [47] to significantly reduce the ADC’s power consumption. Extending our analysis to derive MSE performance guarantees for a modulo ADC with a 1-bit quantizer would therefore be an interesting and valuable direction for future research.
III-C Computational Complexity
We now derive the computational complexities of our proposed recovery algorithms in Sections II-B and II-C. The primary bottleneck of the proposed recovery algorithm for modulo ADC with 1-bit folding information signal is the computation of $\mathbf{V}^{\dagger}_{\mathcal{S}}\in\mathbb{C}^{|\mathcal{S}|\times K}$, where $K$ and $|\mathcal{S}|$ are the number of discrete frequencies in the OOB region and number of folding instances, respectively. For a singular value decomposition (SVD)-based calculation of $\mathbf{V}^{\dagger}_{\mathcal{S}}$, the computational complexity is in $\mathcal{O}\left(K|\mathcal{S}|\cdot\min(K,|\mathcal{S}|)\right)$. Since $|\mathcal{S}|\leq K$ when Lemma 1 holds, the computational complexity of the proposed recovery method in Section II-B is $\mathcal{O}\left(K|\mathcal{S}|^{2}\right)$. On the other hand, the proposed recovery algorithm for modulo ADC without the 1-bit folding information signal still requires the same processing steps as the former in addition to the OMP-based support recovery for estimating $c[n]$. The computational complexity of naive implementation of OMP is in $\mathcal{O}\left(NK+K|\mathcal{S}|^{2}+|\mathcal{S}|^{3}\right)$, where $N$ is the number of samples within an observation window [53]. Since $\mathcal{O}\left(K|\mathcal{S}|^{2}\right)$ is in $\mathcal{O}\left(NK+K|\mathcal{S}|^{2}+|\mathcal{S}|^{3}\right)$ and $\mathcal{O}(|\mathcal{S}|^{3})$ is in $\mathcal{O}(K|\mathcal{S}|^{2})$, the computational complexity of the proposed recovery method in Section II-C is in $\mathcal{O}\left(NK+K|\mathcal{S}|^{2}\right)$.
The need to estimate the 1-bit folding information $c[n]$ incurs some significant computational cost to the recovery procedure, especially when the modulo ADC is operating at high sampling rates. To show this, consider a BL signal $f(t)$ whose 99% of its energy is contained in the time window $(-\frac{T_{w}}{2},\frac{T_{w}}{2})$ for some constant $T_{w}$ (in seconds). Suppose further that $f(t)<\lambda^{\prime}$ outside this time window. The proposed recovery algorithm needs to process the modulo ADC output samples in this time window so that truncation error will be at most 1% of the signal energy. For a modulo ADC with sampling rate $f_{s}$, we have $N=\lceil T_{w}f_{s}\rceil$ samples, $K=\lceil(1-\rho)N\rceil$, and $|\mathcal{S}|\leq K$. Thus, the computational complexity of the proposed OMP-based recovery method scales quadratically with the sampling rate.
The derived computational complexity suggests that the proposed OMP-based recovery could be impractical for high-speed ADCs that need to capture large observation window. Faster sparse recovery algorithms exist such as the iterative soft-thresholding algorithm (ISTA) used in LASSO-$B^{2}R^{2}$ [20]. We emphasize that the primary focus of our current work is to establish MSE performance guarantees in the presence of quantization noise. Our rationale for selecting OMP is its established theoretical guarantees for perfect support recovery under the bounded noise setting.
@@ IV Case Study: Simultaneous Acquisition of Weak and Strong Signals
We now investigate the case of simultaneous acquisition of weak and strong signals that occupy different frequency bands, as shown in Figure 3. The modulo ADC block is based on Fig. 2a and the digital channelizer implements a digital filter that can be tuned to the desired channel. This case is particularly interesting due to the conflicting requirements to capture the two signals. Setting the dynamic range too high to capture the strong signal can bury the weak signal in quantization noise. Meanwhile, setting the dynamic range too low could distort the strong signal. This scenario is also of practical importance. For example, hybrid radar fusion systems, such as the one studied in [54], involve a similar setup, where strong line-of-sight signals and weak echo signals occupy separate frequency bands.
A recent work on USF [37] considers the issue of weak and strong components. However, their mathematical formulation of the problem only assumes a two-tone input signal with different amplitudes. Moreover, performance improvements of modulo ADC over conventional ADCs are only presented via numerical simulations. To the best of our knowledge, a theoretical analysis of the simultaneous acquisition of weak and strong signals using modulo ADC has not been conducted. Our goal in this section is to provide a more general mathematical formulation of the problem and rigorous analysis to justify the advantage of modulo ADC over conventional ADCs in this case.
@@
where $f_{1}(t)$ and $f_{2}(t)$ are bandpass signals. Denote $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ the frequency supports of $f_{1}(t)$ and $f_{2}(t)$, respectively. We impose $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\varnothing$ and $\mathcal{F}_{1}\cup\mathcal{F}_{2}\in(-\omega_{m}/2,+\omega_{m}/2)$, i.e., $f_{1}(t)$ and $f_{2}(t)$ have different frequency supports and they both reside inside $(-\omega_{m}/2,+\omega_{m}/2)$. The bandwidths of $f_{1}(t)$ and $f_{2}(t)$ are $\omega_{1}$ and $\omega_{2}$, respectively. Without loss of generality, we set $\|f_{1}(t)\|_{2}^{2}=1$ and $\|f_{2}(t)\|_{2}^{2}=1$. The simultaneous acquisition of strong and weak signals is modeled by setting $\alpha_{1}\gg\alpha_{2}$. Since $f(t)$ is a bandlimited signal with energy $\alpha_{1}^{2}+\alpha_{2}^{2}$, we have $\|f(t)\|_{\infty}\leq\sqrt{\frac{\omega_{m}(\alpha_{1}^{2}+\alpha_{2}^{2})}{\pi}}$ by [55].
@@
At this point, one might consider separating the two signal components in the analog domain using a tunable bandpass filter. This approach enables us to select which signal to capture before the ADC stage. However, tunable filter with high selectivity are costly – both in terms of hardware resources and design complexity. More precisely, additional analog circuitry is required to vary the center frequency and bandwidth. Tuning elements such as varactors and switched-capacity arrays add complexity, control circuitry, and sensitivity to process, voltage, and temperature (PVT) variations. Furthermore, achieving high selectivity in tunable filters typically requires high-order analog filter designs. However, such high-order filters tend to introduce significant group delay near their cutoff frequencies, leading to signal distortion [56]. In contrast, a tunable digital filter can be implemented by varying the filter coefficients in the registers. Furthermore, achieving a linear phase response is straightforward in digital filters by imposing symmetry on the filter coefficients [56].
@@
Suppose $b>3$, $\mathrm{OF}>3$, and 1-bit folding information is available to the reconstruction procedure. Then the NMSE incurred when recovering the $i$-th component under the modified reconstruction procedure is
@@ V Numerical Results
In this section, we perform the following simulations to substantiate the proposed theory. The aim is to show that, for the same number of quantization bits used for quantization, the modulo ADC surpasses the conventional ADC lacking the modulo operator in terms of quantization noise reduction. This is done by analyzing how the mean squared error (MSE), which represents quantization noise power, changes with the oversampling factor (OF).
Simulation-1: In this simulation, we consider the modulo ADC system with side information as depicted in Fig. 2a. Specifically, we analyze the variation of MSE, as described in Theorem 1 and equation (12), as the OF value varies.
@@
where $\text{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ for any $x\neq 0$. This signal entails a maximal frequency component of $f_{m}=50$ Hz and a temporal duration of $T=1$ second. Random coefficients $A_{i}$ are drawn uniformly from the interval [1, 1][-1,\ 1]. While time offsets $\tau_{i}$ are chosen randomly within the interval $\left(\frac{T}{4},\ \frac{3T}{4}\right)$, this is to ensure that the residual signal is negligible outside the interval (0,T)(0,\ T). Additionally, we normalize $f(t)$ such that $\|f(t)\|_{\infty}=1$. Consequently, the modulo threshold is $\lambda^{\prime}=\frac{1}{\mathrm{OF}-2}$. We consider the number of bits $b$ equal to $4$ and sweep the OF values from $3$ to $15$. The dither signal and uniform quantizer are configured according to these parameters. For each OF value, we compute the MSE between true samples, $f[n]$, and the estimated samples, $\hat{f}[n]$. The experiment is repeated for $20000$ i.i.d. noise realizations and the errors are averaged over all realizations.
@@
Fig. 4 depicts the results obtained with this setting. It is evident from the figure that the MSE of the $b$-modulo ADC with 1-bit folding information decays much faster than that of the conventional $b$-bit ADC without the modulo operator. Therefore, from Fig. 4, we conclude that the quantization noise suppression of modulo ADC is better than that of conventional ADC. Moreover, it is worth noting that for $\text{OF}=3$, $\lambda^{\prime}=\|f(t)\|_{\infty}$, and the MSE values of both modulo ADC and conventional ADC are equal. Fig. 4 also depicts theoretical MSE values obtained using Theorem 1 and equation (12). Note that both theoretical predictions match the simulated results. This justifies the accuracy of our analytical results and further emphasizes the fact that the MSE decays with the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ and $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ for modulo ADC and conventional ADC, respectively. As a final remark, the use of the extra 1-bit side information in Fig. 2a only incurs one additional OR gate to the setup [21]. This penalty is fixed regardless of $b$. From a power efficiency viewpoint, this additional OR gate has minimal impact on the power consumption of modulo ADC.
Simulation-2: In this simulation, we examine the modulo ADC system without the additional 1-bit side information. Specifically, we analyze the MSE variation, as detailed in Theorem 3 and equation (12), across different OF values. The input signal settings align with those of Simulation-1, featuring $b=4$, OF ranging from $4$ to $15$, an OMP stopping criteria parameter of $\eta=0.8$, and $20000$ i.i.d noise realizations. Fig. 5 shows the variation of MSE, between true samples and estimated samples, for this setting as OF varies. The figure demonstrates that modulo ADCs with and without extra-bit side information exhibit superior quantization noise suppression compared to the conventional ADC. Furthermore, our theoretical predictions align well with the simulated results; highlighting that the MSE diminishes at rates of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ and $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ for the modulo ADC and conventional ADC, respectively.
@@
In this simulation, we analyze modulo ADC systems with and without additional 1-bit information, as well as conventional ADCs, to capture both weak and strong signals occupying distinct frequency bands. Specifically, we validate the upper bound on the NMSE for varying OF values, as stated in Corollary 1 and (20), for both strong and weak signal components. The analysis employs the BL signal model described in (18), with parameters $\alpha_{1}=1$ and $\alpha_{2}=0.25$. The signals $f_{1}(t)$ and $f_{2}(t)$ are synthesized using the $\text{sinc}(\cdot)$ function, where the bandwidth of $f_{1}(t)$ is $20\,\mathrm{Hz},20\,\mathrm{Hz}]$, and the bandwidth of $f_{2}(t)$ is $90\,\mathrm{Hz},-50\,\mathrm{Hz}]\cup[50\,\mathrm{Hz},90\,\mathrm{Hz}]$. The simulation parameters are configured as follows: $b=4$, OF ranging from $4$ to $11$, an OMP stopping criteria parameter of $\eta=0.7$, and $5000$ i.i.d. noise realizations. Fig. 6 and Fig. 7 illustrate the variation in NMSE across different OF values for recovering the strong and weak signal components, respectively. These results demonstrate that the modulo ADC outperforms the conventional ADC in the simultaneous acquisition of both strong and weak amplitude components that occupy different frequency bands in a BL signal. Furthermore, these simulation results validate the proposed theoretical upper bound on NMSE. Therefore, the NMSE decreases at rates of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ and $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ for the modulo ADC and conventional ADC, respectively, during the acquisition of strong and weak signal components in a BL signal.
Simulation-4: In the above simulations, we demonstrated that the modulo ADC outperforms the conventional ADC when using both the proposed OMP-based recovery algorithm and and the algorithm with 1-bit folding information. As noted in Section I, several other recovery algorithms for modulo ADCs have been proposed in the literature, including HOD [12], prediction-based [13], and Fourier-domain-based methods [15, 17]. In this simulation, we evaluate and compare the NMSE performance of the modulo ADC using these different recovery algorithms. We adopt the same simulation setup as in Simulation-1, with $\eta=0.8$ and 2000 i.i.d. noise realizations. Fig. 8 plots the NMSE versus OF for $b=4$, comparing the performance of the conventional ADC and the modulo ADC with various recovery algorithms. Notably, the proposed OMP-based and 1-bit-aided recovery methods outperform all other approaches. The relatively poor performance of HOD can be attributed to its sensitivity to noise. Furthermore, we observe that when $b=6$, the performance of the HOD, prediction-based, and B2R2 methods becomes comparable to that of the OMP-based and 1-bit-aided algorithms. These results are presented in Fig. 9. Overall, this analysis highlights the superior performance of the proposed OMP-based and 1-bit-aided recovery techniques, particularly under lower quantizer resolutions.
Figure 8: Numerical NMSE vs. OF for modulo and conventional ADCs with $b=4$. Here, the modulo ADC employs various recovery algorithms, including HOD, prediction-based, Fourier-domain-based (B2R2), the proposed OMP-based method, and with extra 1-bit information.
Figure 9: Numerical NMSE vs. OF for modulo and conventional ADCs with $b=6$. Here, the modulo ADC employs various recovery algorithms, including HOD, prediction-based, Fourier-domain-based (B2R2), the proposed OMP-based method, and with extra 1-bit information.
Simulation-5:
@@
Figure 10: Numerical NMSE results (in dB) for different OF and $b$ values of modulo ADC (with/without extra 1-bit information) and conventional ADC.
In the previous simulations, the OF and $b$ were chosen to satisfy the theoretical conditions for accurate recovery. Here, we examine how the performance of the modulo ADC is affected when these sufficient conditions are not met. Specifically, we compare the NMSE of a classical ADC with that of modulo ADCs, with and without the additional 1-bit information, under various combinations of OF and $b$ that violate the sufficient conditions. The simulation uses the same input signal as in Simulation-1 with $\eta=0.8$, and averages results over 2000 i.i.d. noise realizations. The NMSE results are shown in Fig. 10, leading to the following observations:
  • For $b=2$, the classical ADC outperforms the modulo ADC (with and without 1-bit information) for all OF $\geq 2$.
  • For $b=3$, the classical ADC outperforms the modulo ADC without 1-bit information and performs comparably to the modulo ADC with 1-bit information for all OF $\geq 2$.
@@
Simulation-6: Above simulations use the dithered quantization framework. Here, we demonstrate that even without a dither signal (i.e., using the conventional quantization framework), the modulo ADC combined with the proposed recovery algorithms outperforms the classical ADC. For this analysis, we consider an input signal with the same settings as Simulation-1: $b=4$, an oversampling factor (OF) ranging from $4$ to $11$, an OMP stopping criterion parameter of $\eta=0.8$, a modulo threshold $\lambda=\frac{1}{\mathrm{OF}-2}$, and $5000$ i.i.d. noise realizations. Fig. 11 illustrates the variation in MSE across different OF values for both modulo and classical ADCs. The results show that the modulo ADC, with or without 1-bit information, consistently outperforms the classical ADC.
Figure 11: Numerical MSE vs. OF result for modulo ADC with and without extra 1-bit side information, and conventional ADC.
VI Conclusion